OpenStudy (hannahwaller5):

Will Give Medal! Solve 2x2 + 12x − 14 = 0 by completing the square. 2(x + 6)2 = 26; x = −6 + square root 13, x = −6 − square root 13 2(x + 6)2 = 20; x = −6 + square root 20, x = −6 − square root 20 2(x + 3)2 = 32; x = 1, x = −7 2(x + 3)2 = 32; x = 13, x = 19 So far i got: 2x^2+12x-14=0 2(x^2+12x )-14=0 2(x^2+12x+36)-(14+72)=0 2(x+6)^2-86+86=0+86 2(x+6)^2=86 After i do this (if i did it right) im not sure what to do next...

OpenStudy (janu16):

um starting from beginning 2x^2+12x-14=0 factor out the common term by 2

OpenStudy (hannahwaller5):

Ok so 2(x^2+12x) -14=0

OpenStudy (janu16):

yes but put that -14 in the prathesis

OpenStudy (hannahwaller5):

oh ok

OpenStudy (janu16):

than factor it out by 2

OpenStudy (janu16):

what do you have now

OpenStudy (hannahwaller5):

What do you mean...? Is it supposed to look like this 2(x^2+12x-14)

OpenStudy (janu16):

no so you have 2(x^2+6x-7)=0

OpenStudy (janu16):

12 divideed by 2 is 6

OpenStudy (janu16):

-14 divided by 2 is -7

OpenStudy (hannahwaller5):

Oh sorry i didnt know you were supposed to do that... ok

OpenStudy (janu16):

do you see what im saying?

OpenStudy (hannahwaller5):

Yes now i do

OpenStudy (janu16):

its fine

OpenStudy (janu16):

so you have 2(x^2+6x-7)=0 now Which two numbers add up to 6 and multiply to -7?

OpenStudy (hannahwaller5):

none...

OpenStudy (janu16):

are you sure

OpenStudy (hannahwaller5):

yea you cant multiply anything and get 7...? i think lol

OpenStudy (janu16):

what is -1 plus 7

OpenStudy (janu16):

and what is -1 times 7

OpenStudy (hannahwaller5):

Oh my gosh!! i fell really dumb... ok

OpenStudy (janu16):

do -1 and 7 right?

OpenStudy (janu16):

Rewrite the expression as (x−1)(x+7)

OpenStudy (janu16):

so we have 2(x−1)(x+7)=0

OpenStudy (hannahwaller5):

Well -1 times 7 equals -7 but they dont add up to 6... and ok

OpenStudy (janu16):

umm -1 plus 7 is 6 lol. try doing it again

OpenStudy (hannahwaller5):

Ugh i am like not even paying attention today!!!! sorry i completely forgot the negative

OpenStudy (janu16):

it happens

OpenStudy (janu16):

2(x−1)(x+7)=0 When will (x−1)(x+7) equal zero?

OpenStudy (hannahwaller5):

um... im not completely sure..

OpenStudy (janu16):

ok add when will -1 = to 0

OpenStudy (janu16):

when you add +1 right?

OpenStudy (hannahwaller5):

when you add 1

OpenStudy (janu16):

good when will POSITIVE 7 become 0

OpenStudy (janu16):

@Hannahwaller5

OpenStudy (hannahwaller5):

Sorry i was doing something when you subtract 7

OpenStudy (janu16):

when you subtract 7 Great! x=1,−7 is that one of the choice?

OpenStudy (hannahwaller5):

Yup thank you (:

OpenStudy (janu16):

so we are done! np